Consider the linear operator T:C^3 be defined by
T(z1,z2,z3)= (z1+iz2, iz1-2z2+iz3,-iz2+z3).
Compute T^* and check whether T is self-adjoint.
Check whether T is unitary.
1
Expert's answer
2019-03-22T11:56:52-0400

Given the rules


"z_1\to z_1+iz_2=w_1""z_2\to iz_1-2z_2+iz_3=w_2""z_3\to -iz_2+z_3=w_3"

For linear operator 𝑇 the matrix representation is


"T=\begin{bmatrix}
 1 & i & 0 \\
 i & -2 & i \\ 0 & -i & 1
\end{bmatrix}"

We recall, that an operator T* is called adjoint for the linear operator  T if


"for \ all\ x, y \in \mathbb{C}^3 \ (Tx, y)=(x, T^*y)."

The matrix representation for T* can be found as


"T^*=(\overline{T})^T=\overline{(T^T)}"

where TT denotes the transponse

"\overline{T} \" denotes the matrix with complex conjugated entries.

In our case


"T^*=\begin{bmatrix}
 1 & -i & 0 \\
 -i & -2 & i \\ 0 & -i & 1
\end{bmatrix} \mathrlap{\,/}{=} \ T"

The adjoint operator 


"T^*(z_1, z_2, z_3)=(z_1-iz_2, -iz_1-2z_2+iz_3, -iz_2+z_3)."

Therefore, T is not selfadjoint.


We recall, that a unitary operator is a bounded linear operator on a Hilbert space that satisfies 


"U^*U=UU^*=I_3,"

where U* is the adjoint of U.

In our case


"T^*T=\begin{bmatrix}
 1 & -i & 0 \\
 -i & -2 & i \\ 0 & -i & 1
\end{bmatrix}*\begin{bmatrix}
 1 & i & 0 \\
 i & -2 & i \\ 0 & -i & 1
\end{bmatrix}=\begin{bmatrix}
 2 & 3i & 1 \\
 -3i & 6 & -i \\ 1 & i & 2
\end{bmatrix} \mathrlap{\,/}{=} \ I_3"

Therefore, T is not unitary.